OpenStudy (anonymous):

What is the largest positive integer n such that 1457 and 1754 leave the same remainder when divided by n?

OpenStudy (anonymous):

\[1754=1457 \mod n \rightarrow \left( 1754-1457 \right)=0 \mod n\]

OpenStudy (anonymous):

1754-1457 = 297 so mod n => 297 = 0?

OpenStudy (anonymous):

yes, so what must n be?

OpenStudy (anonymous):

I'm just guessing but 297? Uhgg.. I don't know XD

OpenStudy (anonymous):

btw, they don't have a congreunce symbol so i had to use an = instead. 1457 = k*n + r, 1754 = c*n + r where r is as large as possible. then if we subtract the two, we get (1754 - 1457) = (c-k)*n + (r - r) if we mod n each side, then we get the previous... \[1457\mod297 = 269,\text{ } 1754\mod297 = 269\]

OpenStudy (anonymous):

sorry, n is as large as possible

OpenStudy (anonymous):

Thanks!